OpenStudy (anonymous):

Suppose that $1800 is initially invested in an account at a fixed interest rate, compounded continuously. Suppose also that, after three years, the amount of money in the account is $1999 . Find the interest rate per year. Write your answer as a percentage. Do not round any intermediate computations, and round your percentage to the nearest hundredth.

OpenStudy (anonymous):

N=b * g^t 1999=1800*g^3 g³=1999/1800 g=cbrt(1999/1800)=1.11055556 which is 11% per year.

OpenStudy (anonymous):

nearest hundredth --> 11.06 %

OpenStudy (anonymous):

THANK YOU SO MUCH I COULDNT FIGURE THIS ONE OUT AT ALL...NICE EXPLANATION BTW ;-)